How much material is needed for the right triangular sail on a sailboat with a height of 4.5 m and hypotenuse of

5.5 m? Round your answer to the the nearest tenth of a meter squared.


1
Expert's answer
2020-11-04T16:32:32-0500

The Pythagorean Theorem

If "a"  and "b"  are the lengths of the legs of a right triangle and "c" is the length of the hypotenuse, then the sum of the squares of the lengths of the legs is equal to the square of the length of the hypotenuse


"c^2=a^2+b^2"

Given "a=4.5m, c=5.5 m"

Find the length of the second leg "b"


"b=\sqrt{c^2 -a^2 }"

The area of the right triangle


"Area=\dfrac{1}{2}ab"

"Area=\dfrac{1}{2}a\sqrt{c^2 -a^2 }"

Substitute


"Area=\dfrac{1}{2}(4.5m)\sqrt{(5.5m)^2 -(4.5m)^2 }\approx7.1m^2"

7.1 square meters of material is needed for the right triangular sail on a sailboat with a height of 4.5 m and hypotenuse of 5.5 m.